# Singleton algorithms for the Constraint Satisfaction Problem

### v1

@@ 6.4 Characterization of the singleton BLP+AIP
Suppose $n,\ell\in\mathbb{N}$, $\mathcal{N}$ is a minion whose $n$-ary part is finite, $h\colon\mathcal{M}_{\mathrm{S(BLP+AIP)}}\to\mathcal{N}$ is a minion homomorphism. Then $h(\mathcal{M}_{\mathrm{S(BLP+AIP)}})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric object.
@@
Let us check conditions (1)-(3) of Theorem 6.4. Condition (1) follows from the definition of $\phi$. Let us prove that condition (2) is satisfied for every $\sigma\in\Omega$. For each $i$ appearing in $\sigma$ (viewed as a tuple) let $u(i)\in[n]$ be the position of an overlined block containing only $i$ (such a block exists since $\sigma$ is palette). Then both matrices of $\phi(u(i))^{\sigma}$ have $1$ at the $i$-th position and 0 elsewhere. Hence, $(\phi(u(\sigma(1)))\circ\dots\circ\phi(u(\sigma(N))))^{\sigma}=I^{\sigma}$. Condition (3) holds because adding $I$ makes the matrices skeleton. By Theorem 6.4 we can find $M\in\mathcal{M}_{\mathrm{S(BLP+AIP)}}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for all $(\sigma_{1},\sigma_{2})\in\omega$. Hence, $h(M)$ is an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric object. ∎
@@
Suppose for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots$, $m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette block symmetric function. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
@@
Choose an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette block symmetric function $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, where $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$. For $j\in[S]$ and $i\in[k_{j}]$, let $\alpha_{x}^{j,i}$ be a tuple of length $m_{j,i}$, where each element $a\in A$ appears $s_{m_{j,i}}^{j}(x^{a})$ times. For $j\in\{S+1,S+2,\dots,n\}$ and $i\in[k_{j}]$, let $\alpha_{x}^{j,i}$ be a tuple of length $m_{j,i}$ where each element $a\in A$ appears $s_{m_{j,i}}^{1}(x^{a})$ times. Let us show that a solution to $\mathcal{I}^{\mathbb{B}}$ can be defined by $s(x):=f(\alpha_{x}^{1,1},\dots,\alpha_{x}^{1,k_{1}},\dots,\alpha_{x}^{n,1},\dots,\alpha_{x}^{n,k_{n}})$ for every variable $x$. Let $C=R(x_{1},\dots,x_{k})$ be a constraint of $\mathcal{I}$. For $j\in[S]$ and $i\in[k_{j}]$ let $M_{j,i}$ be the matrix whose columns are tuples from $R$ where each tuple $\mathbf{a}$ appears $s_{m_{j,i}}^{j}(C^{\mathbf{a}})$ times. Similarly, for $j\in\{S+1,S+2,\dots,n\}$ and $i\in[k_{j}]$ let $M_{j,i}$ be the matrix whose columns are tuples from $R$ where each tuple $\mathbf{a}$ appears $s_{m_{j,i}}^{1}(C^{\mathbf{a}})$ times. Our assumptions about $f$ imply that $f(M_{1,1},\dots,M_{1,k_{1}},\dots,M_{n,1},\dots,M_{n,k_{n}})=(s(x_{1}),\dots,s(x_{k}))$. Since $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, it follows that $(s(x_{1}),\dots,s(x_{k}))\in R^{\mathbb{B}}$. Thus, $s$ is a solution of $\mathcal{I}$ in $\mathbb{B}$. ∎
@@
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric function for every $\ell,n\in\mathbb{N}$;
@@
    for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette block symmetric function.

### v2

@@ 5.4 Characterization of the singleton BLP+AIP
Suppose $n,\ell\in\mathbb{N}$, $\mathcal{N}$ is a minion whose $n$-ary part is finite, $h\colon\mathcal{M}_{\mathrm{S(BLP+AIP)}}\to\mathcal{N}$ is a minion homomorphism. Then $h(\mathcal{M}_{\mathrm{S(BLP+AIP)}})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric object.
@@
Let us check conditions (1)-(3) of Theorem 5.4. Condition (1) follows from the definition of $\phi$. Let us prove that condition (2) is satisfied for every $\sigma\in\Omega$. For each $i$ appearing in $\sigma$ (viewed as a tuple) let $u(i)\in[n]$ be the position of an overlined block containing only $i$ (such a block exists since $\sigma$ is palette). Then both matrices of $\phi(u(i))^{\sigma}$ have $1$ at the $i$-th position and 0 elsewhere. Hence, $(\phi(u(\sigma(1)))\circ\dots\circ\phi(u(\sigma(N))))^{\sigma}=I^{\sigma}$. Condition (3) holds because adding $I$ makes the matrices skeleton. By Theorem 5.4 we can find $M\in\mathcal{M}_{\mathrm{S(BLP+AIP)}}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for all $(\sigma_{1},\sigma_{2})\in\omega$. Hence, $h(M)$ is an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric object. ∎
@@
Suppose for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots$, $m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette symmetric function. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
@@
Choose an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette symmetric function $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, where $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$. For $j\in[S]$ and $i\in[k_{j}]$, let $\alpha_{x}^{j,i}$ be a tuple of length $m_{j,i}$, where each element $a\in A$ appears $s_{m_{j,i}}^{j}(x^{a})$ times. For $j\in\{S+1,S+2,\dots,n\}$ and $i\in[k_{j}]$, let $\alpha_{x}^{j,i}$ be a tuple of length $m_{j,i}$ where each element $a\in A$ appears $s_{m_{j,i}}^{1}(x^{a})$ times. Let us show that a solution to $\mathcal{I}^{\mathbb{B}}$ can be defined by $s(x):=f(\alpha_{x}^{1,1},\dots,\alpha_{x}^{1,k_{1}},\dots,\alpha_{x}^{n,1},\dots,\alpha_{x}^{n,k_{n}})$ for every variable $x$. Let $C=R(x_{1},\dots,x_{k})$ be a constraint of $\mathcal{I}$. For $j\in[S]$ and $i\in[k_{j}]$ let $M_{j,i}$ be the matrix whose columns are tuples from $R$ where each tuple $\mathbf{a}$ appears $s_{m_{j,i}}^{j}(C^{\mathbf{a}})$ times. Similarly, for $j\in\{S+1,S+2,\dots,n\}$ and $i\in[k_{j}]$ let $M_{j,i}$ be the matrix whose columns are tuples from $R$ where each tuple $\mathbf{a}$ appears $s_{m_{j,i}}^{1}(C^{\mathbf{a}})$ times. Our assumptions about $f$ imply that $f(M_{1,1},\dots,M_{1,k_{1}},\dots,M_{n,1},\dots,M_{n,k_{n}})=(s(x_{1}),\dots,s(x_{k}))$. Since $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, it follows that $(s(x_{1}),\dots,s(x_{k}))\in R^{\mathbb{B}}$. Thus, $s$ is a solution of $\mathcal{I}$ in $\mathbb{B}$. ∎
@@
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric function for every $\ell,n\in\mathbb{N}$;
@@
    for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette symmetric function.